OpenStudy (anonymous):

A hockey puck leaves a player's stick with a speed of 9.90m/s and slides 33.0m before coming to rest. Find the coefficient of kinetic friction between the puck and the ice.

OpenStudy (anonymous):

Let the weight of the puck be mg. But first you have to find the decelleration of the puck, using the formula v^2=u^2-2as, where v is final speed, u is the initial one, a is acceleration and s is distance, making a subject of the formula you get (0^2-9.9^2)/2*33= -1.485m/s^2. Going to force equation, you F-R=ma, where F is driving force, R is the resistive one and m is mass. Your F is 0 , because it in not given that something is pushing the puck, therefore R=-1.485m N, because you are not given the mass. For this equation, we shall use a different letter from R, because this means contact force. Lets say, instead of our R found, we put F. F=$R, where F is the force, R is contact normal force and $ is the kinetic friction. Replace values and solve fo $. R is mg or 10m. Therefore 1.485m=10m$ and $ is equal to 0.15